OpenStudy (anonymous):

is this right? 15/22t + 8/4t = 15/22t + 2/4t = 37/22t

jhonyy9 (jhonyy9):

(15t+22(2t) ----------- = (15t+44t)/22 = 59t/22 22

OpenStudy (anonymous):

15/22t + 8/4t Are t located in denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

22t and 4t are the denominators

jhonyy9 (jhonyy9):

so from 8/4t how you have got 2/4t ?

OpenStudy (anonymous):

i divided 8/4t by 2

jhonyy9 (jhonyy9):

and 8 divided by 2 is 2 ?

OpenStudy (anonymous):

oh

jhonyy9 (jhonyy9):

so 8/4t you can simplifie by 4 and will get 2/t

OpenStudy (anonymous):

Common denominator is 22t: (15 + 2 * 22)/ 22t = 59/ 22t

jhonyy9 (jhonyy9):

fave is ok ?

jhonyy9 (jhonyy9):

is it ok ?

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

this all

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is this correct? z^2-81/z divide by z+9/z+5 I got z^2-45/z as the answer

jhonyy9 (jhonyy9):

ok than this is ok so you can give medal for ... to ...

jhonyy9 (jhonyy9):

(z-9)(z+9) z+5 (z-9)(z+5) ---------- * ------ = ----------- z z+9 z

OpenStudy (anonymous):

z^2-81/z divide by z+9/z+5 (z^2-81)/z * (z +5)/ (z +9) (z -9)(z+9)/z * (z +5)/ (z +9) => ( z -9 ) * (z + 5) / z

jhonyy9 (jhonyy9):

z^2 -4z -45 ----------- z